# Impact of 3D macro physics and nuclear physics on the $p$ nuclei in $\mathrm{O}$-$\mathrm{C}$ shell mergers

### v1

@@ Appendix A Correlations of nuclear reaction rates
A caveat of $r_{\mathrm{P}}$ for this method of varying the reaction rates is that it not distinguish between the photo-disintegration and corresponding capture rate because the same variation factor is applied to both. All correlated rates are reported according to their photo-disintegration rates, but as shown by the upper left plot of Figure A1 for $\mathrm{{}^{74}Se}$ and $\mathrm{{}^{75}Se}(\gamma,n)\mathrm{{}^{74}Se}$ this results in a production term having a negative correlation because $\mathrm{{}^{74}Se}(n,\gamma)\mathrm{{}^{75}Se}$ is also modified in the same way. As explained in Section 3.1, the reactions in this shell are not balanced and both a destruction and production term could be relevant at different locations, although as Figure 8 shows for heavier species the $(\gamma,n)$ rate is typically much stronger.
@@ Appendix B Results of varying the ingestion rate
Figure B1: The overproduction compared to initial of the $p$ nuclei for the 3D-inspired mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.58\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
Figure B2: The overproduction compared to initial of the $p$ nuclei for the 3×3$\times$3 ×3D-inspired mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.64\;\mathrm{\,\mathrm{dex}}$ excluding the no ingestion case. $\mathrm{OP}=0$ is the initial amount.
Figure B3: The overproduction compared to initial of the $p$ nuclei for the 10×10$\times$10 ×3D-inspired mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. Arrows denote $\mathrm{OP}$ out of bounds and the true $\mathrm{OP}$ is written above. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.78\;\mathrm{\,\mathrm{dex}}$ excluding the no ingestion case. $\mathrm{OP}=0$ is the initial amount.
Figure B4: The overproduction compared to initial of the $p$ nuclei for the 50×50$\times$50 ×3D-inspired mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. Arrows denote $\mathrm{OP}$ out of bounds and the true $\mathrm{OP}$ is written above. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.84\;\mathrm{\,\mathrm{dex}}$ excluding the no ingestion case. $\mathrm{OP}=0$ is the initial amount.
@@ Appendix C Results of varying the input nuclear reactions
Figure C1: Histogram showing the spread due to varying $(\gamma,p)$, $(\gamma,n)$, $(\gamma,\alpha)$ and corresponding capture rates for unstable $n$-deficient isotopes from $\mathrm{Se}$$\mathrm{Po}$ for the 3D-inspired mixing scenario. Colour and size both correspond to the logarthimic binning of Monte Carlo runs. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.59\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
Figure C2: Histogram showing the spread due to varying $(\gamma,p)$, $(\gamma,n)$, $(\gamma,\alpha)$ and corresponding capture rates for unstable $n$-deficient isotopes from $\mathrm{Se}$$\mathrm{Po}$ for the 3×3$\times$3 ×3D-inspired mixing scenario. Colour and size both correspond to the logarthimic binning of Monte Carlo runs. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.69\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
Figure C3: Histogram showing the spread due to varying $(\gamma,p)$, $(\gamma,n)$, $(\gamma,\alpha)$ and corresponding capture rates for unstable $n$-deficient isotopes from $\mathrm{Se}$$\mathrm{Po}$ for the 10×10$\times$10 ×3D-inspired mixing scenario. Colour and size both correspond to the logarthimic binning of Monte Carlo runs. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.76\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
Figure C4: Histogram showing the spread due to varying $(\gamma,p)$, $(\gamma,n)$, $(\gamma,\alpha)$ and corresponding capture rates for unstable $n$-deficient isotopes from $\mathrm{Se}$$\mathrm{Po}$ for the 50×50$\times$50 ×3D-inspired mixing scenario. Colour and size both correspond to the logarthimic binning of Monte Carlo runs. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.79\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
@@
${}^{74}\mathrm{Se}$ ${}^{75}\mathrm{Se}(\gamma,n)$ $-0.93$ $-0.18$ ${}^{144}\mathrm{Sm}$ ${}^{142}\mathrm{Sm}(\gamma,n)$ $-0.19$ $-0.02$
${}^{78}\mathrm{Kr}$ ${}^{79}\mathrm{Kr}(\gamma,n)$ $-0.88$ $-0.28$ ${}^{142}\mathrm{Sm}(\gamma,p)$ $-0.17$ $-0.02$
${}^{84}\mathrm{Sr}$ ${}^{85}\mathrm{Sr}(\gamma,n)$ $-0.88$ $-0.28$ ${}^{143}\mathrm{Sm}(\gamma,n)$ $-0.25$ $-0.03$
${}^{92}\mathrm{Mo}$ ${}^{93}\mathrm{Mo}(\gamma,n)$ $-0.94$ $-0.07$ ${}^{146}\mathrm{Sm}(\gamma,n)$ $0.20$ $0.03$
@@
${}^{94}\mathrm{Mo}$ ${}^{93}\mathrm{Mo}(\gamma,n)$ $0.97$ $0.06$ ${}^{150}\mathrm{Gd}(\gamma,\alpha)$ $-0.15$ $-0.02$
${}^{96}\mathrm{Ru}$ ${}^{97}\mathrm{Ru}(\gamma,n)$ $-0.88$ $-0.12$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.47$ $0.06$
@@
${}^{98}\mathrm{Ru}$ ${}^{97}\mathrm{Ru}(\gamma,n)$ $0.36$ $0.02$ ${}^{152}\mathrm{Gd}$ ${}^{152}\mathrm{Dy}(\gamma,\alpha)$ $-0.40$ $-0.01$
${}^{100}\mathrm{Pd}(\gamma,p)$ $0.62$ $0.04$ ${}^{154}\mathrm{Dy}(\gamma,\alpha)$ $-0.15$ $-0.00$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.51$ $-0.03$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $0.39$ $0.00$
${}^{102}\mathrm{Pd}$ ${}^{100}\mathrm{Pd}(\gamma,p)$ $-0.29$ $-0.05$ ${}^{156}\mathrm{Dy}$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.18$ $-0.06$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.30$ $-0.05$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $0.74$ $0.26$
${}^{103}\mathrm{Pd}(\gamma,n)$ $-0.66$ $-0.12$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.18$ $0.05$
${}^{106}\mathrm{Cd}$ ${}^{107}\mathrm{Cd}(\gamma,n)$ $-0.86$ $-0.16$ ${}^{158}\mathrm{Dy}$ ${}^{158}\mathrm{Er}(\gamma,\alpha)$ $-0.23$ $-0.01$
@@
${}^{109}\mathrm{Cd}(\gamma,n)$ $-0.46$ $-0.06$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.18$ $0.00$
${}^{110}\mathrm{Sn}(\gamma,p)$ $0.16$ $0.02$ ${}^{162}\mathrm{Er}$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.18$ $-0.06$
${}^{113}\mathrm{In}$ ${}^{113}\mathrm{Sn}(\gamma,n)$ $0.91$ $0.37$ ${}^{160}\mathrm{Er}(\gamma,n)$ $-0.18$ $-0.06$
${}^{112}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.17$ $-0.03$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $-0.26$ $-0.07$
${}^{113}\mathrm{Sn}(\gamma,n)$ $-0.83$ $-0.15$ ${}^{161}\mathrm{Er}(\gamma,n)$ $0.21$ $0.06$
${}^{114}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.15$ $-0.01$ ${}^{166}\mathrm{Yb}(\gamma,\alpha)$ $0.53$ $0.14$
@@
${}^{122}\mathrm{Xe}(\gamma,n)$ $-0.18$ $-0.01$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.17$ $0.05$
${}^{122}\mathrm{Xe}(\gamma,p)$ $0.20$ $0.02$ ${}^{164}\mathrm{Er}$ ${}^{164}\mathrm{Yb}(\gamma,n)$ $-0.24$ $-0.09$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $0.41$ $0.04$ ${}^{164}\mathrm{Yb}(\gamma,\alpha)$ $-0.58$ $-0.32$
@@
${}^{122}\mathrm{Xe}(\gamma,p)$ $0.17$ $0.01$ ${}^{168}\mathrm{Yb}$ ${}^{168}\mathrm{Hf}(\gamma,\alpha)$ $-0.28$ $-0.14$
@@
${}^{120}\mathrm{Te}$ ${}^{121}\mathrm{Te}(\gamma,n)$ $-0.71$ $-0.09$ ${}^{176}\mathrm{W}(\gamma,\alpha)$ $0.20$ $0.07$
@@
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.32$ $-0.04$ ${}^{174}\mathrm{Hf}$ ${}^{174}\mathrm{W}(\gamma,n)$ $-0.21$ $-0.03$
${}^{124}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,n)$ $-0.17$ $-0.04$ ${}^{174}\mathrm{W}(\gamma,\alpha)$ $-0.40$ $-0.08$
${}^{122}\mathrm{Xe}(\gamma,p)$ $-0.24$ $-0.06$ ${}^{178}\mathrm{W}(\gamma,n)$ $-0.15$ $-0.03$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.45$ $-0.15$ ${}^{178}\mathrm{W}(\gamma,\alpha)$ $0.42$ $0.07$
${}^{123}\mathrm{Xe}(\gamma,n)$ $0.16$ $0.03$ ${}^{180}\mathrm{Ta}$ ${}^{179}\mathrm{Ta}(\gamma,n)$ $-0.91$ $-0.02$
${}^{125}\mathrm{Xe}(\gamma,n)$ $-0.46$ $-0.16$ ${}^{180}\mathrm{W}$ ${}^{180}\mathrm{Os}(\gamma,n)$ $-0.28$ $-0.07$
${}^{126}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.34$ $-0.09$ ${}^{180}\mathrm{Os}(\gamma,\alpha)$ $-0.52$ $-0.21$
@@
${}^{127}\mathrm{Xe}(\gamma,n)$ $-0.25$ $-0.07$ ${}^{184}\mathrm{Os}$ ${}^{185}\mathrm{Os}(\gamma,n)$ $0.33$ $0.02$
${}^{126}\mathrm{Ba}(\gamma,p)$ $-0.18$ $-0.04$ ${}^{184}\mathrm{Pt}(\gamma,\alpha)$ $-0.44$ $-0.03$
${}^{126}\mathrm{Ba}(\gamma,\alpha)$ $-0.32$ $-0.09$ ${}^{186}\mathrm{Pt}(\gamma,n)$ $0.17$ $0.01$
${}^{130}\mathrm{Ba}$ ${}^{126}\mathrm{Ba}(\gamma,\alpha)$ $-0.17$ $-0.01$ ${}^{186}\mathrm{Pt}(\gamma,\alpha)$ $-0.16$ $-0.01$
${}^{128}\mathrm{Ba}(\gamma,n)$ $-0.19$ $-0.01$ ${}^{188}\mathrm{Pt}(\gamma,n)$ $-0.17$ $-0.01$
@@
${}^{131}\mathrm{Ba}(\gamma,n)$ $-0.79$ $-0.05$ ${}^{190}\mathrm{Pt}$ ${}^{190}\mathrm{Hg}(\gamma,n)$ $-0.28$ $-0.03$
${}^{132}\mathrm{Ba}$ ${}^{131}\mathrm{Ba}(\gamma,n)$ $0.66$ $0.09$ ${}^{190}\mathrm{Hg}(\gamma,\alpha)$ $-0.51$ $-0.06$
${}^{133}\mathrm{Ba}(\gamma,n)$ $-0.61$ $-0.08$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.02$
${}^{138}\mathrm{La}$ ${}^{137}\mathrm{La}(\gamma,n)$ $-0.71$ $-0.37$ ${}^{196}\mathrm{Hg}$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $-0.71$ $-0.41$
${}^{136}\mathrm{Ce}$ ${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.38$ $-0.05$ ${}^{197}\mathrm{Pb}(\gamma,n)$ $-0.16$ $-0.06$
@@
${}^{138}\mathrm{Nd}(\gamma,\alpha)$ $-0.16$ $-0.02$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.30$ $0.14$
@@
${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.30$ $-0.01$
${}^{138}\mathrm{Nd}(\gamma,\alpha)$ $-0.22$ $-0.01$

### v2

@@ Appendix A Correlations of nuclear reaction rates
A caveat of $r_{\mathrm{P}}$ for this method of varying the reaction rates is that it not distinguish between the photo-disintegration and corresponding capture rate because the same variation factor is applied to both. All correlated rates are reported according to their photo-disintegration rates, but as shown by the upper left plot of Figure A1 for $\mathrm{{}^{74}Se}$ and $\mathrm{{}^{75}Se}(\gamma,n)\mathrm{{}^{74}Se}$ this results in a production term having a negative correlation because $\mathrm{{}^{74}Se}(n,\gamma)\mathrm{{}^{75}Se}$ is also modified in the same way. As explained in Section III.1, the reactions in this shell are not balanced and both a destruction and production term could be relevant at different locations, although as Figure 8 shows for heavier species the $(\gamma,n)$ rate is typically much stronger.
@@ Appendix B Results of varying the ingestion rate
Figure B1: The overproduction compared to initial of the $p$ nuclei for the 3D-inspired mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.58\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount. Average $\mathrm{OP}$ for each scenario are provided as dashed lines and corresponds to the values presented in Table 1.
Figure B2: The overproduction compared to initial of the $p$ nuclei for the $3\times$3D-inspired mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.64\;\mathrm{\,\mathrm{dex}}$ excluding the no ingestion case. $\mathrm{OP}=0$ is the initial amount. Average $\mathrm{OP}$ for each scenario are provided as dashed lines and corresponds to the values presented in Table 1.
Figure B3: The overproduction compared to initial of the $p$ nuclei for the $10\times$3D-inspired mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. Arrows denote $\mathrm{OP}$ out of bounds and the true $\mathrm{OP}$ is written above. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.78\;\mathrm{\,\mathrm{dex}}$ excluding the no ingestion case. $\mathrm{OP}=0$ is the initial amount. Average $\mathrm{OP}$ for each scenario are provided as dashed lines and corresponds to the values presented in Table 1.
Figure B4: The overproduction compared to initial of the $p$ nuclei for the $50\times$3D-inspired mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. Arrows denote $\mathrm{OP}$ out of bounds and the true $\mathrm{OP}$ is written above. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.84\;\mathrm{\,\mathrm{dex}}$ excluding the no ingestion case. $\mathrm{OP}=0$ is the initial amount. Average $\mathrm{OP}$ for each scenario excluding “No Merger” are provided as dashed lines and corresponds to the values presented in Table 1.
@@ Appendix C Results of varying the input nuclear reactions
Figure C1: Histogram showing the spread due to varying $(\gamma,p)$, $(\gamma,n)$, $(\gamma,\alpha)$ and corresponding capture rates for unstable $n$-deficient isotopes from $\mathrm{Se}$$\mathrm{Po}$ for the 3D-inspired mixing scenario. Colour and size both correspond to the logarithmic binning of Monte Carlo runs. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.59\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
Figure C2: Histogram showing the spread due to varying $(\gamma,p)$, $(\gamma,n)$, $(\gamma,\alpha)$ and corresponding capture rates for unstable $n$-deficient isotopes from $\mathrm{Se}$$\mathrm{Po}$ for the $3\times$3D-inspired mixing scenario. Colour and size both correspond to the logarithmic binning of Monte Carlo runs. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.69\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
Figure C3: Histogram showing the spread due to varying $(\gamma,p)$, $(\gamma,n)$, $(\gamma,\alpha)$ and corresponding capture rates for unstable $n$-deficient isotopes from $\mathrm{Se}$$\mathrm{Po}$ for the $10\times$3D-inspired mixing scenario. Colour and size both correspond to the logarithmic binning of Monte Carlo runs. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.76\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
Figure C4: Histogram showing the spread due to varying $(\gamma,p)$, $(\gamma,n)$, $(\gamma,\alpha)$ and corresponding capture rates for unstable $n$-deficient isotopes from $\mathrm{Se}$$\mathrm{Po}$ for the $50\times$3D-inspired mixing scenario. Colour and size both correspond to the logarithmic binning of Monte Carlo runs. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.79\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
@@
$^{74}\mathrm{Se}$ $^{75}\mathrm{Se}(\gamma,n)$ 0.93-0.93 0.18-0.18 $^{144}\mathrm{Sm}$ $^{142}\mathrm{Sm}(\gamma,n)$ 0.19-0.19 0.02-0.02
$^{78}\mathrm{Kr}$ $^{79}\mathrm{Kr}(\gamma,n)$ 0.88-0.88 0.28-0.28 $^{142}\mathrm{Sm}(\gamma,p)$ 0.17-0.17 0.02-0.02
$^{84}\mathrm{Sr}$ $^{85}\mathrm{Sr}(\gamma,n)$ 0.88-0.88 0.28-0.28 $^{143}\mathrm{Sm}(\gamma,n)$ 0.25-0.25 0.03-0.03
$^{92}\mathrm{Mo}$ $^{93}\mathrm{Mo}(\gamma,n)$ 0.94-0.94 0.07-0.07 $^{146}\mathrm{Sm}(\gamma,n)$ $0.20$ $0.03$
@@
$^{94}\mathrm{Mo}$ $^{93}\mathrm{Mo}(\gamma,n)$ $0.97$ $0.06$ $^{150}\mathrm{Gd}(\gamma,\alpha)$ 0.15-0.15 0.02-0.02
$^{96}\mathrm{Ru}$ $^{97}\mathrm{Ru}(\gamma,n)$ 0.88-0.88 0.12-0.12 $^{196}\mathrm{Pb}(\gamma,n)$ $0.47$ $0.06$
@@
$^{98}\mathrm{Ru}$ $^{97}\mathrm{Ru}(\gamma,n)$ $0.36$ $0.02$ $^{152}\mathrm{Gd}$ $^{152}\mathrm{Dy}(\gamma,\alpha)$ 0.40-0.40 0.01-0.01
$^{100}\mathrm{Pd}(\gamma,p)$ $0.62$ $0.04$ $^{154}\mathrm{Dy}(\gamma,\alpha)$ 0.15-0.15 0.00-0.00
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.51-0.51 0.03-0.03 $^{160}\mathrm{Er}(\gamma,\alpha)$ $0.39$ $0.00$
$^{102}\mathrm{Pd}$ $^{100}\mathrm{Pd}(\gamma,p)$ 0.29-0.29 0.05-0.05 $^{156}\mathrm{Dy}$ $^{159}\mathrm{Er}(\gamma,n)$ 0.18-0.18 0.06-0.06
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.30-0.30 0.05-0.05 $^{160}\mathrm{Er}(\gamma,\alpha)$ $0.74$ $0.26$
$^{103}\mathrm{Pd}(\gamma,n)$ 0.66-0.66 0.12-0.12 $^{202}\mathrm{Pb}(\gamma,n)$ $0.18$ $0.05$
$^{106}\mathrm{Cd}$ $^{107}\mathrm{Cd}(\gamma,n)$ 0.86-0.86 0.16-0.16 $^{158}\mathrm{Dy}$ $^{158}\mathrm{Er}(\gamma,\alpha)$ 0.23-0.23 0.01-0.01
@@
$^{109}\mathrm{Cd}(\gamma,n)$ 0.46-0.46 0.06-0.06 $^{202}\mathrm{Pb}(\gamma,n)$ $0.18$ $0.00$
$^{110}\mathrm{Sn}(\gamma,p)$ $0.16$ $0.02$ $^{162}\mathrm{Er}$ $^{159}\mathrm{Er}(\gamma,n)$ 0.18-0.18 0.06-0.06
$^{113}\mathrm{In}$ $^{113}\mathrm{Sn}(\gamma,n)$ $0.91$ $0.37$ $^{160}\mathrm{Er}(\gamma,n)$ 0.18-0.18 0.06-0.06
$^{112}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.17-0.17 0.03-0.03 $^{160}\mathrm{Er}(\gamma,\alpha)$ 0.26-0.26 0.07-0.07
$^{113}\mathrm{Sn}(\gamma,n)$ 0.83-0.83 0.15-0.15 $^{161}\mathrm{Er}(\gamma,n)$ $0.21$ $0.06$
$^{114}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.15-0.15 0.01-0.01 $^{166}\mathrm{Yb}(\gamma,\alpha)$ $0.53$ $0.14$
@@
$^{122}\mathrm{Xe}(\gamma,n)$ 0.18-0.18 0.01-0.01 $^{202}\mathrm{Pb}(\gamma,n)$ $0.17$ $0.05$
$^{122}\mathrm{Xe}(\gamma,p)$ $0.20$ $0.02$ $^{164}\mathrm{Er}$ $^{164}\mathrm{Yb}(\gamma,n)$ 0.24-0.24 0.09-0.09
$^{122}\mathrm{Xe}(\gamma,\alpha)$ $0.41$ $0.04$ $^{164}\mathrm{Yb}(\gamma,\alpha)$ 0.58-0.58 0.32-0.32
@@
$^{122}\mathrm{Xe}(\gamma,p)$ $0.17$ $0.01$ $^{168}\mathrm{Yb}$ $^{168}\mathrm{Hf}(\gamma,\alpha)$ 0.28-0.28 0.14-0.14
@@
$^{120}\mathrm{Te}$ $^{121}\mathrm{Te}(\gamma,n)$ 0.71-0.71 0.09-0.09 $^{176}\mathrm{W}(\gamma,\alpha)$ $0.20$ $0.07$
@@
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.32-0.32 0.04-0.04 $^{174}\mathrm{Hf}$ $^{174}\mathrm{W}(\gamma,n)$ 0.21-0.21 0.03-0.03
$^{124}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,n)$ 0.17-0.17 0.04-0.04 $^{174}\mathrm{W}(\gamma,\alpha)$ 0.40-0.40 0.08-0.08
$^{122}\mathrm{Xe}(\gamma,p)$ 0.24-0.24 0.06-0.06 $^{178}\mathrm{W}(\gamma,n)$ 0.15-0.15 0.03-0.03
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.45-0.45 0.15-0.15 $^{178}\mathrm{W}(\gamma,\alpha)$ $0.42$ $0.07$
$^{123}\mathrm{Xe}(\gamma,n)$ $0.16$ $0.03$ $^{180}\mathrm{Ta}$ $^{179}\mathrm{Ta}(\gamma,n)$ 0.91-0.91 0.02-0.02
$^{125}\mathrm{Xe}(\gamma,n)$ 0.46-0.46 0.16-0.16 $^{180}\mathrm{W}$ $^{180}\mathrm{Os}(\gamma,n)$ 0.28-0.28 0.07-0.07
$^{126}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.34-0.34 0.09-0.09 $^{180}\mathrm{Os}(\gamma,\alpha)$ 0.52-0.52 0.21-0.21
@@
$^{127}\mathrm{Xe}(\gamma,n)$ 0.25-0.25 0.07-0.07 $^{184}\mathrm{Os}$ $^{185}\mathrm{Os}(\gamma,n)$ $0.33$ $0.02$
$^{126}\mathrm{Ba}(\gamma,p)$ 0.18-0.18 0.04-0.04 $^{184}\mathrm{Pt}(\gamma,\alpha)$ 0.44-0.44 0.03-0.03
$^{126}\mathrm{Ba}(\gamma,\alpha)$ 0.32-0.32 0.09-0.09 $^{186}\mathrm{Pt}(\gamma,n)$ $0.17$ $0.01$
$^{130}\mathrm{Ba}$ $^{126}\mathrm{Ba}(\gamma,\alpha)$ 0.17-0.17 0.01-0.01 $^{186}\mathrm{Pt}(\gamma,\alpha)$ 0.16-0.16 0.01-0.01
$^{128}\mathrm{Ba}(\gamma,n)$ 0.19-0.19 0.01-0.01 $^{188}\mathrm{Pt}(\gamma,n)$ 0.17-0.17 0.01-0.01
@@
$^{131}\mathrm{Ba}(\gamma,n)$ 0.79-0.79 0.05-0.05 $^{190}\mathrm{Pt}$ $^{190}\mathrm{Hg}(\gamma,n)$ 0.28-0.28 0.03-0.03
$^{132}\mathrm{Ba}$ $^{131}\mathrm{Ba}(\gamma,n)$ $0.66$ $0.09$ $^{190}\mathrm{Hg}(\gamma,\alpha)$ 0.51-0.51 0.06-0.06
$^{133}\mathrm{Ba}(\gamma,n)$ 0.61-0.61 0.08-0.08 $^{196}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.02$
$^{138}\mathrm{La}$ $^{137}\mathrm{La}(\gamma,n)$ 0.71-0.71 0.37-0.37 $^{196}\mathrm{Hg}$ $^{196}\mathrm{Pb}(\gamma,n)$ 0.71-0.71 0.41-0.41
$^{136}\mathrm{Ce}$ $^{138}\mathrm{Nd}(\gamma,n)$ 0.38-0.38 0.05-0.05 $^{197}\mathrm{Pb}(\gamma,n)$ 0.16-0.16 0.06-0.06
@@
$^{138}\mathrm{Nd}(\gamma,\alpha)$ 0.16-0.16 0.02-0.02 $^{202}\mathrm{Pb}(\gamma,n)$ $0.30$ $0.14$
@@
$^{138}\mathrm{Nd}(\gamma,n)$ 0.30-0.30 0.01-0.01
$^{138}\mathrm{Nd}(\gamma,\alpha)$ 0.22-0.22 0.01-0.01